y= 1/4x - 3; (-8, ) , ( , 1)
Thanks in advanced!

Answer by macston(5194)

y= 1/4x - 3; (-8, ) , ( , 1)
First ordered pair:
x=-8
y=(1/4)(-8)-3
y=-2-3=-5
y=-5
ANSWER 1: (-8,-5)
Second ordered pair:
y=1
1=(1/4)x-3 Add 3 to each side.
4=(1/4)x Multiply each side by 4.
16=x
ANSWER 2: (16,1)
y=1/4x-3; (-8,-5) , (16,1)